OpenStudy (anonymous):

simplify. -3b^4 + b^3 - (b^3 + b^4 - 3)

OpenStudy (ajprincess):

Combine like terms. Can u do?

OpenStudy (anonymous):

i'm confused

OpenStudy (ajprincess):

-3b^4 + b^3 - (b^3 + b^4 - 3) =-3b^4+b^3-b^3-b^4+3 What is -3b^4-b^4=?

OpenStudy (anonymous):

-81b i don't know what b^4 is

OpenStudy (ajprincess):

b^4 is term with the power 4. b^4=\(b^4\) -3-1=?

OpenStudy (anonymous):

-4

OpenStudy (ajprincess):

-3b^4-b^4=-4b^4 Do u get nw?

OpenStudy (ajprincess):

The terms with the same power are called like terms. U can add or subtract only like terms. do u follow me?

OpenStudy (anonymous):

i think so

OpenStudy (ajprincess):

b^3-b^3=?

OpenStudy (anonymous):

0

OpenStudy (ajprincess):

u r right. Nw -3b^4+b^3-b^3-b^4+3=?

OpenStudy (anonymous):

-81

OpenStudy (anonymous):

-81 - 256b

OpenStudy (ajprincess):

hmm no.

OpenStudy (anonymous):

ok im lost

OpenStudy (ajprincess):